OpenStudy (anonymous):

Can someone help/show me how to do these: 1. 1/6(12-6m)=5m-6 2. x/4+5+x=15 4. a/3+2a/5=1/2 5. x+7/9-x-5/6=5/3

OpenStudy (beginnersmind):

Can you do these? a) 3x=18 b) 4x+x=15 c) 3u=12-u

OpenStudy (anonymous):

a: x=6 b: x=3 c: x=3

OpenStudy (beginnersmind):

Ok, let's look at the third one. y/2-y/3=11 grouping the terms with y together  then multiply both sides by 6 (cont)